OpenStudy (anonymous):

The sum of three integers is 92. The second number is three times the first number. The third number is ten less than twice the first number. Find the integers.

TheSmartOne (thesmartone):

a, b, and c are our 3 integers. Can you make equations using a, b, and c using what you were given in your question?

OpenStudy (anonymous):

A + B + C = 92

OpenStudy (anonymous):

B= A(3)

OpenStudy (anonymous):

C= (A-10) * 2

TheSmartOne (thesmartone):

Now substitue the 2 last equations into the first one :)

OpenStudy (anonymous):

A + (A(3)) + ((A - 10)(2)) = 92

OpenStudy (anonymous):

i think the equation for c is close, but not correct

OpenStudy (anonymous):

c = 2a - 10 i think

OpenStudy (anonymous):

I am not sure

OpenStudy (anonymous):

but yeah after you fix that and plug it in, you just solve for A

TheSmartOne (thesmartone):

Yes, c = 2a -10 is correct. Solve for a, then you can solve for b and c by using the last 2 equations.